# Can AI Dream of Unseen Galaxies? Conditional Diffusion Model for Galaxy Morphology Augmentation

### v1

@@ Appendix D Additional Sanity Check on our D-ETG Samples
We here present the r-band magnitude and stellar mass distribution to complete our discussions in Section 4.3. As we have We present the r-band magnitude and stellar mass distributions to complement our discussions in Section 4.3. Having identified 520 new D-ETGs using ML models enhanced with synthetic images, we compare their properties with the D-ETG sample from Kaviraj et al. (2012) to validate their shared physical consistency and with a control sample of early-type galaxies (ETGs) to ensure rigorous sample control.
@@
Figure D4: Cumulative distribution functions (CDFs) comparing r-band magnitudes and stellar masses of our newly identified D-ETGs (red), the D-ETG sample from Kaviraj et al. (2012) (blue), and control ETGs (black). The close alignment between our sample and Kaviraj et al. (2012) confirms consistent physical properties for dusty early-type galaxies, while systematic differences from control ETGs validate the distinctiveness of the dusty morphology subset and the rigor of our sample control.
Appendix E Compared to GPT-4o
@@
As we have demonstrated that our fine-tuned diffusion model can generate high-quality galaxy images adhere to instruct we compare our diffusion model’s results with the state-of-the-art industrial text-to-image model GPT-4o [^6] https://openai.com/index/introducing-4o-image-generation/ here, having demonstrated in the main text that our model generates high-fidelity, scientifically meaningful galaxy images. As shown in Figure E5, industrial text-to-image models effectively capture key visual features and morphological concepts of galaxy images but require further refinement to accurately replicate the systematic characteristics of real observational data.
@@
For future astrophysical AI development, leveraging pre-trained industrial generative models as a foundation is more pragmatic. As illustrated in our study, starting from Stable-Diffusion-v1-5 and training on only 27,910 high-quality annotated galaxy images yielded strong performance. Industrial models, having been trained on billions of high-quality image-text pairs, already encode the majority of general visual concepts, streamlining downstream domain adaptation for specialized astronomical applications.
@@
Figure E5: Comparison of galaxy images from SDSS, our model’s simulations, and GPT-4o. Annotated capital letters correspond to generation prompts listed in Table 1. While GPT-4o has encode those visual concepts related to galaxy morphology, it may not fully reproduce real observed image characteristics, e.g., noise level and point spread functions. As demonstrated in our study, starting from industrial pre-trained models (e.g., Stable-Diffusion-v1-5) and performing domain adaptation on science-specific datasets efficiently bridges this gap, enabling accurate reproduction of observational features critical for astrophysical applications.

### v2

@@ Appendix E Additional Sanity Check on our D-ETG Samples
We here present the r-band magnitude and stellar mass distribution to complete our discussions in Section IV.3. Having identified 520 new D-ETGs using ML models enhanced with synthetic images, we compare their properties with the D-ETG sample from Kaviraj et al. (2012) to validate their shared physical consistency and with a control sample of early-type galaxies (ETGs) to ensure rigorous sample control.
@@
Figure E6: Cumulative distribution functions (CDFs) comparing r-band magnitudes and stellar masses of our newly identified D-ETGs (red), the D-ETG sample from Kaviraj et al. (2012) (blue), and control ETGs (black). The close alignment between our sample and Kaviraj et al. (2012) confirms consistent physical properties for dusty early-type galaxies, while systematic differences from control ETGs validate the distinctiveness of the dusty morphology subset and the rigor of our sample control.
Appendix F Compared to GPT-4o
@@
As we have demonstrated that our fine-tuned diffusion model can generate high-quality galaxy images adhere to instruct we compare our diffusion model’s results with the state-of-the-art industrial text-to-image model GPT-4o [^8] https://openai.com/index/introducing-4o-image-generation/ here, having demonstrated in the main text that our model generates high-fidelity, scientifically meaningful galaxy images. As shown in Figure F8, industrial text-to-image models effectively capture key visual features and morphological concepts of galaxy images but require further refinement to accurately replicate the systematic characteristics of real observational data.
To more comprehensively compare images generated by our GalaxySD with those from GPT-4o, we conducted evaluations using the metrics introduced in the paper, namely realism and consistency. Because GPT-4o’s image generation function is usage-limited, each user could submit only a small number of text-to-image requests, making it infeasible to compute diversity in a statistically meaningful way. Consequently, our comparison with GPT-4o focuses primarily on realism and consistency, while the diversity metric is reported only for our model to highlight its ability to produce varied outputs under identical prompts. Figure F7 indicates the differences, our scores minus GPT-4o’s scores, on both axes. Using differences rather than absolute scores could directly visualize the performance gap and makes it clear whether our model performs better (positive values) or worse (negative values) than GPT-4o. Nearly all cases lie above or to the right of the gray dashed zero lines, with a few located directly on the lines, indicating that our GalaxySD generally outperforms GPT-4o in both realism and consistency.
Figure F7: Scatter plot of the differences in realism and consistency scores between galaxy images generated by our GalaxySD and GPT-4o. The horizontal and vertical axes represent the differences, i.e., our scores minus GPT-4o’s scores. The error bars indicate ±10% of the absolute value of each difference, reflecting the uncertainty or variability in the measurements due to the limited number of GPT-4o images. Different colors and markers distinguish the detailed morphological prompts.
@@
For future astrophysical AI development, leveraging pretrained industrial generative models as a foundation is more pragmatic. As illustrated in our study, starting from Stable-Diffusion-v1-5 and training on only 27,910 high-quality annotated galaxy images yielded strong performance. Industrial models, having been trained on billions of high-quality image-text pairs, already encode the majority of general visual concepts, streamlining downstream domain adaptation for specialized astronomical applications.
@@
Figure F8: Comparison of galaxy images from SDSS, our model’s simulations, and GPT-4o. Annotated capital letters correspond to generation prompts listed in Table 1. As demonstrated in our study, starting from industrial pretrained models (e.g., Stable-Diffusion-v1-5) and performing domain adaptation on science-specific datasets efficiently bridges the gap between general visual concepts and galaxy-specific observational features, enabling accurate reproduction of observational features critical for astrophysical applications.